OpenStudy (anonymous):

3(x - 6) + 2 = 4(x + 2) - 21 solve for x

mathslover (mathslover):

\[\huge{3x-18+2=4x+8-21}\] \[\huge{3x-4x-18-8+2+21=0}\] \[\huge{-x-26+2+21=0}\] \[\huge{-x-3=0}\] \[\huge{-x=3}\] \[\huge{x=-3}\]

OpenStudy (anonymous):

so the anwser is -3?

mathslover (mathslover):

yes @tnicole22

OpenStudy (anonymous):

thank you!